OpenStudy (anonymous):

How heavy do passengers feel as they enter the loop of a roller coaster that begins with a 90m hill followed by a loop 40m in diameter?

OpenStudy (valpey):

The equation for centripetal acceleration is: \[a_r = \frac{v^2}{r}\] velocity at the entrance to the loop will be given by\[ v=\sqrt{2gh}\] "Heavyness" is basically "how many Gs" so \[\frac{g+a_r}{g}\]

OpenStudy (anonymous):

Ok, so... v=sqrt(2*9.8*90) centripetal acceleration=v^2/20? Weight= [9.8+(a sub r)]/9.8?

OpenStudy (valpey):

I think sa. 10 Gs sound high, but I also don't think this is a real-world roller coaster.

OpenStudy (valpey):

*I think so.

OpenStudy (anonymous):

How would I calculate the gravity? I was just going off 9.8 m/s^2

OpenStudy (valpey):

Yeah, that is a good approximation. Doesn't matter as it cancels out anyway.

OpenStudy (anonymous):

Didn't even think of that!

OpenStudy (anonymous):

Ok, so working it out I got that the weight would be 90. Would that be correct?

OpenStudy (valpey):

I get \[\frac{1 G+9 Gs}{1}=10Gs\] This has some of the derivations and whatnot: http://www.real-world-physics-problems.com/roller-coaster-physics.html

OpenStudy (anonymous):

How did you get 9G's? \[a _{r}=\frac{ 2*90*G}{ 20}= \frac{ 180G}{ 20}=90G\] Plugging in: \[\frac{ G+90G}{ G}=91?\]

OpenStudy (anonymous):

@mathmate

OpenStudy (mathmate):

I must get some data straight. A roller coaster usually goes up a hill (90m) and then drop down to a loop. If we assume that's the case, then the passenger only falls through 40m (diameter of the loop, or the loop begins 50 m down?

OpenStudy (mathmate):

So let us assume that "entering" the loop means it is at the lowest point where the potential energy is lowest.

OpenStudy (mathmate):

And that it would be 90m below the hill.

OpenStudy (mathmate):

Kinetic energy=(1/2)mv^2=mgh so as Valpey puts it, v=sqrt(2gh)=sqrt(180g). acceleration at bottom is additive to g, so the acceleration (downwards) is v^2/r +g =180g/(40/2) + g = (9+1)g =10g as @Valpey had it.

OpenStudy (mathmate):

@mathsss you probably got one zero too many in 180g/20=9g

OpenStudy (anonymous):

Wasn't thinking! Thank you guys!